# Background Semantics Matter: Cross-Task Feature Exchange Network for Clustered Infrared Small Target Detection With Sky-Annotated Dataset

### v1

@@ III-B Target Synthesis and Composition
The dataset leverages the existing SIRSTv2, creating a broad library of infrared small targets. Each image is meticulously designed to emulate the densely populated surveillance scenarios. Specifically, one to three “dense areas” are defined randomly within the less cluttered sky regions of each base image. These areas, each measuring $20\times 20$ pixels, are populated with 8 to 12 targets selected randomly from the target library. The selected targets are resized to dimensions within $5\times 5$ pixels to maintain scale variability and pasted with a spacing of 1 to 2 pixels between them to simulate realistic clustering.
@@
$T_{new}=T_{back}+B_{box}\times\lambda\times G\ ,\lambda\in[0.5,1]$ (1)
Here, $T_{new}$ denotes the adjusted target appearance, while $T_{back}$ is the original background at the target’s location. Different from mixup, its weight is set to $1$ for enhancing the reality of the synthesized target. $B_{box}$ represents the original target appearance. The scalar $\lambda$, which adjusts the target’s brightness, is randomized to diversify the brightness and contrast of targets, thus enriching the sample space.
@@
The relative offsets $\rho_{x}$ and $\rho_{y}$ determine the Gaussian distribution center along the $x$ and $y$ axes, randomly selected within the range $[0,0.2]$. These values, when multiplied by the target bounding box’s width $w$ and height $h$, yield the actual pixel offsets, ensuring appropriate scaling with the target size. The Gaussian distribution’s spread is controlled by the standard deviations $\sigma_{x}$ and $\sigma_{y}$ along the $x$ and $y$ axes, randomized within $[0.3,0.6]$. This randomness guarantees varying degrees of edge blurring among the simulated targets, reinforcing their realistic appearance.
To further enhance variability, each target undergoes a rotational transformation. The Gaussian matrix’s rotation $\theta$ is applied as:
@@
Here, $\theta$ is randomly selected from the interval $\left[-90^{\circ},90^{\circ}\right]$.
@@ III-C Dataset Description and Characteristics
III-C1 Composition of the Dataset
The DenseSIRST dataset comprises 1024 infrared images, with a total of 13,655 densely clustered small targets. The images in the DenseSIRST dataset cover a wide range of realistic scenarios, including urban areas, mountainous regions, maritime environments, and cloudy scenes. This diversity enables the development and assessment of infrared small target detection algorithms that can generalize well to various real-world applications. The division of training set, validation set, and test set is aligned with the SIRSTv2 dataset.
@@ III-C2 Annotations and Labels
III-C3 Statistical Characteristics of the Dataset
The DenseSIRST dataset is meticulously curated to provide a comprehensive benchmark for clustered infrared small target detection. This part offers a detailed analysis of the dataset’s key characteristics, including the distribution of image sizes, target sizes, local contrast, and brightness levels.
Figure 4: Figure (a) is the image resolution characteristic, the larger the range of circles in the scatter plot, the greater the number of images at that resolution. Figure (b) is the target size characteristic, and the larger the range of circles in the scatter plot, the greater the number of small targets of that size. Figure (c) shows the percentage of different local contrasts for small objects in the dataset, with the lowest local contrast being 0.15. Figure (d) shows the percentage of brightness distribution for small targets in the dataset, and shows that the brightness distribution is wide from dark targets to light targets.
Image Size Distribution: The dataset, as depicted in Fig. 4 (a), includes images with resolutions ranging from $320\times 256$ to $1280\times 1024$ pixels, covering various typical infrared camera resolutions encountered in real-world scenarios. This diversity is crucial for developing detection algorithms that can generalize across different resolutions.
Target Size Distribution: Most targets, as indicated in Fig. 4 (b), are extremely small, predominantly below $5\times 5$ pixels. The $3\times 3$ pixels size is the most frequent, aligning with the characteristics of small target detection in infrared imagery. With a total of 13,655 annotated targets, the dataset provides a rich set of examples for training and evaluation.
Local Contrast Distribution: The vast majority (90%) of targets have local contrast values below $2$, as shown in Fig. 4 (c). This low contrast is indicative of the inherent difficulty in detecting such targets against cluttered backgrounds. Moreover, the presence of nearby targets within dense clusters can inadvertently be included in the background calculation, leading to an underestimation of local contrast and further complicating detection.
Target Brightness Distribution: Fig. 4 (d) reveals that only a minority of targets are the brightest points in their respective images. This characteristic emphasizes the inadequacy of brightness as a sole distinguishing feature for target detection in infrared images.
@@ IV-A Overall Architecture and Training Pipeline
The proposed BAFE-Net adopts a fully convolutional one-stage object detector (FCOS) [34] as its host network. FCOS is a simple yet effective anchor-free object detection framework that has demonstrated superior performance in various computer vision tasks. By leveraging the inherent advantages of FCOS, such as its ability to handle objects of different scales and aspect ratios without the need for predefined anchor boxes, BAFE-Net aims to address the challenges of detecting small, densely clustered targets in infrared imagery.
@@
The overall architecture of BAFE-Net is illustrated in Fig. 5. In addition to the original detection head present in the baseline FCOS, BAFE-Net introduces a segmentation head within the BAFE-Head module. The primary purpose of incorporating this segmentation head is to enable explicit modeling of background semantics, which is crucial for enhancing the network’s ability to distinguish between real targets and background clutter in infrared scenes. By jointly optimizing the detection and segmentation tasks, BAFE-Net learns to capture both object-level and pixel-level contextual information, ultimately leading to improved detection performance.
@@
In addition to the architecture modifications, BAFE-Net incorporates a novel training methodology termed Background-Aware Gaussian Copy-Paste (BAG-CP). This technique is used during the training phase to augment the training dataset with realistically integrated synthetic targets. BAG-CP strategically places synthetic infrared small targets into real scene backgrounds by considering semantic context, which significantly enhances the training process. The use of Gaussian smoothing at the edges of the pasted targets ensures seamless integration, avoiding artificial boundary artifacts that could mislead the learning algorithm.
This context-aware synthesis and integration approach allows BAFE-Net to train on a dataset that closely mimics operational scenarios, where small targets often appear against complex and cluttered backgrounds. By training the network to recognize and handle these challenging conditions effectively, BAFE-Net sets a new standard for performance in infrared small target detection.
IV-B Cross-Task Feature Exchange Module
During the feature extraction process, partial feature exchange is performed between segmentation and classification tasks within this network. The segmentation and classification tasks have different focuses when learning features: segmentation emphasizes local details and contextual information, whereas classification emphasizes global features. Despite their differing objectives, there is a potential synergistic relationship between segmentation and classification tasks. By exchanging features, information can be shared between the two tasks, promoting collaborative learning and enhancing the feature representation capabilities of each task. Segmentation can assist classification in better understanding the structure and semantics of the image, while classification can help segmentation in distinguishing boundaries and details of different categories. Specifically, segmentation features can leverage global information from classification features to improve overall understanding, while classification features can utilize local detail information from segmentation features to enhance precision. Mutual learning between segmentation and classification features can reduce the risk of overfitting and improve the model’s performance in practical applications. Feature exchange enables the model to perform more robustly when facing different tasks. It allows the model to better adapt to various changes and uncertainties, thereby enhancing its generalization ability across different scenarios.
Feature exchange has been proven feasible for several reasons: 1) Complementarity of features: Although segmentation and classification features are different, they are extracted from the same input image and thus possess a degree of complementarity. By exchanging features, each task can utilize useful information from the other task, thereby improving overall performance. 2) Compatibility of features: Deep learning models use continuous and differentiable operations, such as convolution and pooling, when extracting features. Through feature exchange, these operations can still ensure the coherence and differentiability of features, thus not disrupting the training process of the model. 3) Support from existing research: Some studies have already shown that in multi-task learning, sharing or exchanging features can improve the overall performance of tasks. For example, in the Changer [35] paper, channel and spatial exchanges were used and proven feasible.
Feature exchange primarily consists of two methods: channel exchange and spatial exchange. Channel exchange involves directly swapping features at the channel level between different feature maps, while spatial exchange is achieved by swapping features along the spatial dimensions of the feature maps. In the Cross-Task Feature Exchange Module, channel exchange is employed, and experiments have shown that channel exchange performs better than spatial exchange. The experimental results will be presented in Section V.
@@ IV-B Cross-Task Feature Exchange Module
Assume ${x}_{1}$ and ${x}_{2}$ are the input tensors, each with a shape of $(N,C,H,W)$, where $N$ is the batch size, $C$ is the number of channels, and $H$ and $W$ are the height and width, respectively. Let $p$ be a parameter controlling the proportion of channel exchange. Define an exchange mask $E$ with a shape of $C$,
${E}[i]=\begin{cases}1&\text{, if }i\mod\left(\frac{1}{p}\right)=0\\ 0\end{cases}$ (4)
@@
where $i$ is the channel index. Expand the exchange mask $E$ to the shape of $(N,C,H,W)$,
${E}_{\text{expanded}}={E}\otimes{1}_{N\times H\times W}$ (5)
@@
where $\otimes$ denotes the expansion operation, and ${1}_{N\times H\times W}$ is a tensor of ones with a shape of $(N,H,W)$. Define the output tensors ${y}_{1}$ and ${y}_{2}$ ,
${y}_{1}={x}_{1}\odot(1-{E}_{\text{expanded}})+{x}_{2}\odot{E}_{\text{expanded}}$ (6)
${y}_{2}={x}_{2}\odot(1-{E}_{\text{expanded}})+{x}_{1}\odot{E}_{\text{expanded}}$ (7)
where $\odot$ denotes element-wise multiplication.
However, the features to be exchanged are cross-modal. Therefore, an Adapter is introduced to process these features before the exchange. Here, an MLP (Multi-Layer Perceptron) is used as the Adapter. The features to be exchanged are fed into the Adapter separately, and after processing, the features are exchanged. This approach addresses the cross-modal issue between the features.
@@ V Experiments
    Ablation Study on Background-Aware Gaussian Copy-Paste. We compare the performance of various copy-paste strategies to demonstrate the superiority of our BAG-CP method.
@@
    Effectiveness of Background-Aware Feature Exchange. We evaluate the performance of BAFE-Net by comparing different configurations to highlights the contributions of each component in BAFE-Net.
@@ V-A Experimental Settings
2) Implementation Details: Our network was trained on an NVIDIA RTX 3090 GPU platform with 24GB of memory. In our experiments, we employed the DeepInfrared toolkit for object detection. Considering the image sizes in the dataset and computational overhead, for methods within the DeepInfrared toolkit, we set the range of the long edge of training and testing images to [512, 512] while maintaining aspect ratio to avoid image distortion. We choose FocalLoss as the loss function of the classification branch, IoULoss as the loss function of the regression branch, and CrossEntropyLoss as the loss function of the segmentation branch. We choose DAdaptAdam as the optimizer, lr is set to 1.0, and weight decay is set to 0.05. Considering training and convergence time, we trained for 20 epochs within the DeepInfrared framework. Through experimentation, we observed that different parameter settings affect the detection efficacy of dense small objects. Thus, while ensuring basic fairness, we adjusted the parameters of different methods to obtain optimal detection results.
3) Evaluation Metrics:We used mean Average Precision (mAP) with IoU thresholds of 0.5 ($mAP_{07}$) and 0.75 ($mAP_{12}$), and Recall as evaluation metrics for our experiment. They are defined as:
$\text{mAP}=\frac{1}{N}\sum_{i=1}^{N}\text{AP}(i),$ (8)
[…]
$\text{Recall}=\frac{TP}{TP+FN},$ (10)
Here, $N$ is the total number of samples, $TP$, $FP$, and $FN$ denote the number of true positive, false positive, and false negative instances, respectively. Precision $p(r)$ is defined as the proportion of true positive detections out of the total number of positive detections at a given recall level $r$.
TABLE II: Quantitative ablation study evaluates the impact of the Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting within the data augmentation module of BAFE-Net

### v2

@@ III-B Target Synthesis and Composition
The dataset leverages our SIRST V2 dataset, creating a broad library of infrared small targets. Each image is meticulously designed to emulate the densely populated surveillance scenarios. Specifically, one to three “dense areas” are defined randomly within the less cluttered sky regions of each base image. These areas, each measuring $20\times 20$ pixels, are populated with 8 to 12 targets selected randomly from the target library. The selected targets are resized to dimensions within $5\times 5$ pixels to maintain scale variability and pasted with a spacing of 1 to 2 pixels between them to simulate realistic clustering.
@@
$T_{new}=T_{back}+B_{box}\times\lambda\times G\ ,\lambda\in[0.5,1],$ (1)
where $T_{new}$ denotes the adjusted target appearance, while $T_{back}$ is the original background at the target’s location. Different from mixup, its weight is set to $1$ for enhancing the reality of the synthesized target. $B_{box}$ represents the original target appearance. The scalar $\lambda$, which adjusts the target’s brightness, is randomized to diversify the brightness and contrast of targets, thus enriching the sample space.
@@
where the relative offsets $\rho_{x}$ and $\rho_{y}$ determine the Gaussian distribution center along the $x$ and $y$ axes, randomly selected within the range $[0,0.2]$. These values, when multiplied by the target bounding box’s width $w$ and height $h$, yield the actual pixel offsets, ensuring appropriate scaling with the target size. The Gaussian distribution’s spread is controlled by the standard deviations $\sigma_{x}$ and $\sigma_{y}$ along the $x$ and $y$ axes, randomized within $[0.3,0.6]$. This randomness guarantees varying degrees of edge blurring among the simulated targets, reinforcing their realistic appearance. To further enhance variability, each target undergoes a rotational transformation. The Gaussian matrix’s rotation $\theta$ is applied as:
@@
where $\theta$ is randomly selected from the interval $\left[-90^{\circ},90^{\circ}\right]$.
@@ III-C Dataset Description and Characteristics
III-C1 Dataset Composition
The DenseSIRST dataset comprises 1024 infrared images, with a total of 13,655 densely clustered small targets. The images in the DenseSIRST dataset cover a wide range of realistic scenarios, including urban areas, mountainous regions, maritime environments, and cloudy scenes. This diversity enables the development and assessment of detection algorithms that can generalize well to various real-world applications. The division of training, validation, and test set is aligned with the SIRST V2 dataset.
@@ III-C2 Annotations and Labels
III-C3 Statistical Characteristics
The DenseSIRST dataset is meticulously curated to provide a comprehensive benchmark for clustered infrared small target detection. This part offers an analysis of the dataset’s key characteristics.
Target Size Distribution: Most targets, as indicated in Fig. 3 (b), are extremely small, predominantly below $5\times 5$ pixels. The $3\times 3$ pixels size is the most frequent, aligning with the characteristics of small target detection in infrared imagery.
Local Contrast Distribution: The 90% targets have local contrast below $2$, as shown in Fig. 3 (c). This low contrast is indicative of the inherent difficulty in detecting such targets against cluttered backgrounds. Moreover, the presence of nearby targets within dense clusters can inadvertently be included in the background calculation, leading to an underestimation of local contrast and further complicating detection.
Target Brightness Distribution: Fig. 3 (d) reveals that only a minority of targets are the brightest points in their respective images. This characteristic emphasizes the inadequacy of brightness as a sole distinguishing feature for target detection.
@@ IV-A Overall Architecture and Training Pipeline
The overall architecture of BAFE-Net, which adopts the fully convolutional one-stage object detector (FCOS) [37] as its host network, is illustrated in Fig. 4. In addition to the original detection head present in the baseline, BAFE-Net introduces a segmentation head within the BAFE-Head module. The primary purpose of incorporating this segmentation head is to enable explicit modeling of background semantics, which is crucial for enhancing the network’s ability to distinguish between real targets and background clutter in infrared scenes. By jointly optimizing the detection and segmentation tasks, BAFE-Net learns to capture both object-level and pixel-level contextual information. Furthermore, BAFE-Net introduces a dynamic cross-task feature hard-exchange mechanism to facilitate the bidirectional flow of information between the detection and segmentation branches. This dynamic exchange allows the network to adaptively leverage the most relevant features from each task, leading to improved detection performance.
@@
To address the challenge of limited target samples, BAFE-Net incorporates a novel training methodology termed Background-Aware Gaussian Copy-Paste (BAG-CP). This technique strategically places synthetic infrared small targets into real scene backgrounds by considering semantic context, ensuring that the pasted targets are only placed in sky regions. The use of Gaussian smoothing at the edges of the pasted targets ensures seamless integration, avoiding artificial boundary artifacts that could mislead the learning algorithm. By augmenting the training dataset with realistically integrated synthetic targets, BAG-CP allows BAFE-Net to train on a dataset that closely mimics operational scenarios, where small targets often appear against complex and cluttered backgrounds. This context-aware synthesis and integration approach enables BAFE-Net to effectively recognize and handle these challenging conditions, ultimately setting a new standard for performance in infrared small target detection.
IV-B Dynamic Cross-Task Feature Exchange Head
For infrared small target detection, the scarcity of intrinsic target features and the abundance of similar background distractors pose significant challenges. To address this, we propose the Background-Aware Feature Exchange Network (BAFE-Net), which transforms the detection paradigm from a single task focused on the foreground to a multi-task architecture that jointly performs target detection and background semantic segmentation. The motivation behind this cross-task feature exchange is twofold:
  1. 1.
    Complementary Nature of Tasks: Although target detection and background segmentation have different objectives, they share a common goal of understanding the scene. By exchanging features between these tasks, each task can benefit from the complementary information learned by the other, leading to a more comprehensive understanding of the image.
  2. 2.
    Synergistic Relationship: Target detection focuses on identifying the presence and location of small targets, while background segmentation emphasizes the contextual information and local details of the scene. By sharing features between these tasks, target detection can leverage the semantic understanding of the background to better distinguish targets from distractors, while background segmentation can utilize the localized target information to refine its segmentation boundaries.
IV-B1 Dynamic Channel Selection
To further optimize cross-task feature exchange, we propose a dynamic selection mechanism that determines the most relevant feature channels for exchange based on their importance to task performance. The importance of each channel is estimated using an attention mechanism, which assigns weights to channels according to their contribution.
Given an input feature map $\mathbf{X}\in\mathbb{R}^{C\times H\times W}$, where $C$, $H$, and $W$ denote the number of channels, height, and width of the feature map, respectively, we first compute a channel-wise attention weight vector $\mathbf{m}\in\mathbb{R}^{C}$ by a mechanism inspired by the design of our previous work Large Selective Kernel Network (LSKNet) [38]. This is achieved by applying a series of operations to $\mathbf{X}$, including local feature extraction, channel-wise statistics computation, and weighted fusion. The process can be summarized as follows:
$\mathbf{m}$ $=\sigma\left(\mathbf{X}\odot\left(\mathcal{F}(\mathbf{X})\odot\sigma\left(\mathcal{S}\left(\text{avg}(\mathbf{X})\oplus\max(\mathbf{X})\right)\right)\right)\right),$ (4)
where $\odot$ denotes element-wise multiplication, $\oplus$ denotes the concatenation operation, and $\sigma$ is the Sigmoid function. $\max$ and avg are the maximum and average pooled spatial feature descriptors. $\mathcal{S}$ is a convolution operation that reduces the dimensionality of concatenated features. $\mathcal{F}$ is also a convolution operation applied to $\mathbf{X}$ , extracting local features.$\mathbf{m}\in\mathbb{R}^{C}$ is a one-dimensional tensor, and $C$ is the number of channels. Each element $\mathbf{m}$ represents the attention logits for the $i$-th channel.
For each sample, the top $K$ channels with the highest weights are selected, where $K$ is determined as a fraction $p$ of the total number of channels $C$:
$K$ $=C\times p,$ (5)
${V}_{\text{topk}},{I}_{\text{topk}}$ $=\mathrm{TopK}({\mathbf{m}},K),$
where ${V}_{\text{topk}}\in\mathbb{R}^{K}$ represents the top $K$ largest weight values for each sample. ${I}_{\text{topk}}\in\mathbb{R}^{K}$ represents the corresponding channel indices for these weights. This process dynamically selects the channels for exchange, improving the effectiveness of cross-task feature sharing.
IV-B2 Cross-Task Feature Hard-Exchange
Within BAFE-Net, we introduce a cross-task feature hard-exchange mechanism to embed target and background semantics between the two tasks. Feature hard-exchange allows for the direct transfer of learned features from one task to another. By swapping entire feature channels, the receiving task has access to the explicit representation learned by the other task, facilitating a more direct integration of complementary information.
@@ IV-B2 Cross-Task Feature Hard-Exchange
To perform feature exchange, let $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ be the input tensors from the target detection and background segmentation tasks, respectively, each with a shape of $(C,H,W)$, where $C$ is the number of channels, and $H$ and $W$ are the height and width. We define an exchange mask $\mathbf{e}\in\mathbb{R}^{C}$:
$\mathbf{e}[i]=\begin{cases}1,&\mathrm{if}\ i\in I_{\text{topk}}\\ 0,&\mathrm{otherwise}\end{cases},$ (6)
@@
where $i$ represents the channel index. The exchange mask $\mathbf{e}$ is then expanded to the shape of $(C,H,W)$:
$\mathbf{E}_{d}=\mathbf{e}\otimes\mathbf{1}_{H\times W},$ (7)
@@
where $\otimes$ denotes the expansion operation, and $\mathbf{1}_{H\times W}$ is a tensor of ones with a shape of $(H,W)$. $\mathbf{X}_{1}$ and $\mathbf{X}_{2}$ can obtain $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ respectively through Eq. (4) to Eq. (7). The output tensors $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ after the hard feature exchange are computed as:
$\mathbf{Y}_{1}=\mathbf{X}_{1}\odot(\mathbf{1}-\mathbf{E}_{d1})+\mathcal{A}\left(\mathbf{X}_{2}\odot\mathbf{E}_{d2}\right),$ (8)
$\mathbf{Y}_{2}=\mathbf{X}_{2}\odot(\mathbf{1}-\mathbf{E}_{d2})+\mathcal{A}\left(\mathbf{X}_{1}\odot\mathbf{E}_{d1}\right),$ (9)
where $\odot$ denotes element-wise multiplication, $\mathbf{Y}_{1}$ and $\mathbf{Y}_{2}$ represent the output features, and $\mathbf{E}_{d1}$ and $\mathbf{E}_{d2}$ are the dynamic masks. To address the cross-modal nature of features, we introduce an adapter module $\mathcal{A}$, implemented as a multi-layer perceptron (MLP), which projects features into a common space before exchange. This design enables effective cross-modal interaction while preserving modality-specific characteristics.
@@ V Experiments
    Ablation Study on BAG-CP. We compare the performance of various copy-paste strategies to demonstrate the superiority of our BAG-CP method.
@@
    Effectiveness of Background-Aware Feature Exchange. We evaluate the performance of BAFE-Net by comparing different configurations to highlight the contributions of each component in BAFE-Net.
@@ V-A Experimental Settings
2) Implementation Details: Considering the image sizes in the dataset and computational overhead, for methods within the DeepInfrared toolkit, we set the range of the long edge of training and testing images to [512, 512] while maintaining aspect ratio to avoid image distortion. We choose Focal Loss as the loss function of the classification branch, IoU Loss as the loss function of the regression branch, and Cross Entropy Loss as the loss function of the segmentation branch. We choose DAdaptAdam as the optimizer, learning rate is set to 1.0, and weight decay is set to 0.05. Considering training and convergence time, we trained for 20 epochs within the DeepInfrared framework.
3) Evaluation Metrics:We used mean Average Precision (mAP) with IoU thresholds of 0.5 ($mAP_{07}$) and 0.75 ($mAP_{12}$), and Recall as evaluation metrics for our experiment.
TABLE II: Comparison with Other State-of-the-art methods on DenseSIRST.
Method Backbone DenseSIRST
FLOPs $\downarrow$ Params $\downarrow$ mAP07 $\uparrow$ recall07 $\uparrow$ mAP12 $\uparrow$ recall12 $\uparrow$
              One-stage
SSD [39] 87.552G 23.746M 0.211 0.421 0.178 0.424
RetinaNet [40] ResNet50 52.203G 36.330M 0.114 0.510 0.086 0.523
YOLOv3 [41] 50.002G 61.949M 0.233 0.424 0.207 0.413
CenterNet [42] ResNet50 50.278G 32.111M 0.138 0.316 0.124 0.317
FCOS [37] ResNet50 50.291G 32.113M 0.232 0.315 0.204 0.324
ATSS [43] ResNet50 51.504G 32.113M 0.248 0.327 0.202 0.326
CentripetalNet [44] 0.491T 0.206G 0.244 0.259 0.201 0.244
AutoAssign [45] ResNet50 50.555G 36.244M 0.255 0.354 0.180 0.314
GFL [46] ResNet50 52.296G 32.258M 0.264 0.367 0.230 0.317
PAA [47] ResNet50 51.504G 32.113M 0.255 0.545 0.228 0.551
VFNet [48] ResNet50 48.317G 32.709M 0.253 0.336 0.214 0.336
PVT-T [49] 41.623G 21.325M 0.109 0.481 0.093 0.501
YOLOF [50] ResNet50 25.076G 42.339M 0.091 0.009 0.002 0.009
YOLOX [51] 8.578G 8.968M 0.210 0.341 0.180 0.331
TOOD [12] ResNet50 50.456G 32.018M 0.256 0.355 0.226 0.342
DyHead [52] ResNet50 27.866G 38.890M 0.249 0.335 0.189 0.328
DDOD [53] ResNet50 46.514G 32.378M 0.253 0.335 0.230 0.351
RTMDet [54] 51.278G 52.316M 0.229 0.349 0.212 0.350
EfficientDet [55] 34.686G 18.320M 0.146 0.464 0.094 0.517
Two-stage
Faster R-CNN [56] ResNet50 0.759T 33.035M 0.091 0.022 0.015 0.029
Cascade R-CNN [57] ResNet50 90.978G 69.152M 0.136 0.188 0.139 0.194
Grid R-CNN [58] ResNet50 0.177T 64.467M 0.156 0.122 0.104 0.190
Libra R-CNN [59] ResNet50 63.990G 41.611M 0.141 0.142 0.085 0.120
TridentNet [60] ResNet50 0.759T 33.035M 0.091 0.009 0.014 0.021
SABL [61] ResNet50 0.125T 42.213M 0.124 0.104 0.104 0.171
Dynamic R-CNN [62] ResNet50 63.179G 41.348M 0.184 0.235 0.111 0.190
End2End
DETR [63] ResNet50 24.940G 41.555M 0.000 0.000 0.000 0.000
Sparse R-CNN [64] ResNet50 45.274G 0.106G 0.183 0.572 0.154 0.614
Deformable DETR [65] ResNet50 51.772G 40.099M 0.024 0.016 0.018 0.197
Conditional DETR [66] ResNet50 27.143G 40.297M 0.000 0.000 0.000 0.001
DAB-DETR [67] ResNet50 28.939G 43.702M 0.005 0.054 0.000 0.001
$\star$ BAFE-Net (Ours) ResNet18 57.654G 22.31M 0.283 0.335 0.233 0.325
$\star$ BAFE-Net (Ours) ResNet50 71.639G 35.626M 0.274 0.342 0.248 0.338
V-B Comparison with State-of-the-Arts
We benchmarked our BAFE-Net against current leading or state-of-the-art methods on the proposed DenseSIRST dataset, using metrics such as $mAP_{07}$, $mAP_{12}$, and Recall. The experimental results are presented in Table II.
TABLE III: Quantitative ablation study evaluates the impact of the Copy-Paste technique, clustering of pasted objects, Gaussian module, and Background-Aware pasting within the data augmentation module of BAFE-Net